the mean weight of baggage of 25 passengers from USA bound for the Philippines is 31.8 kilos with a variance of 16. test the claim that the population mean baggage weight is equal to 30 kilos at 0.02 level of significance.


1
Expert's answer
2021-09-23T16:40:00-0400

Null hypothesis "H_0:\mu=30."

Alternative hypothesis "H_a:\mu\ne30."

Test statistic: "z=\frac{31.8-30}{\sqrt{\frac{16}{25}}}=2.25."

P-value: "p=2P(Z>2.25)=0.0244."

Since the p-value is greater than 0.02, fail to reject the null hypothesis.

There is no significant evidence that the population mean baggage weight is not equal to 30 kilos.